2. A coin was flipped 80 times and came up tails 54 times. At the 0.05 level of significance, is the coin biased toward tails? Show your decision rule and calculations.
Answer by stanbon(48568)

A coin was flipped 80 times and came up tails 54 times. At the 0.05 level of significance, is the coin biased toward tails? Show your decision rule and calculations.
---------------------------------
Ho: p = 1/2
Ha: p > 1/2
---------------
sample proportion: 54/80 = 0.675
---
test statistic: z(0.675) = (0.675-0.5)/sqrt((1/2)(1/2)/80) = 3.1305
--------------------
Critical value for right-tail test with alpha=5%: z=-1.96 or z=1.96
------------------------------
p-value = 2(P(z>1.96)) = 2(0.000873) = 0.0017
--------------------------------------------------
Decision Rule: If p-value is less than 5%, reject Ho.
---------------------------------------------------
Conclusion: Since p-value is less than 5%, reject Ho ; the cois is
biased toward heads.
==========================================================
Cheers,
Stan H.
-----------------------